OpenStudy (anonymous):

please solve for Y: x+3y=13 2x-y=-9

OpenStudy (anonymous):

Let say: x+3y=13---- No.1 2x-y=-9 -----No.2 From No.2 y=2x+9---- No.3 Sub y=2x+9 into No.1 x+3(2x+9)=13 x+6x+27=13 7x=13-27 =-14 x=-14/7 =-2 Sub x=-2 into No.3 y=2(-2)+9 =9-4 =5:) Try it out! Good luck!